OpenStudy (anonymous):

determine which ratio forms a proportion with 24/28 by writing the ratios in simplest form

OpenStudy (anonymous):

\[\huge \frac{24}{28} = \frac{6 \times 4}{7 \times 4} =\frac{6}{7} \] hence \[\huge \frac{6}{7} would- be-in- proportion- with \frac{24}{28} \]